OpenStudy (anonymous):

Three multiplied by the sum of 4 and a number is the same as 18 more than the number. Find the number. If n is "the number," which equation could be used to solve for the number?

Parth (parthkohli):

\[3(n + 4) = n + 18\]Wanna solve that equation.

OpenStudy (anonymous):

3 ( 4 + n) = 18 + n 12 + 3n = 18 + n 2n - 6 = 0

OpenStudy (anonymous):

2n - 6 = 0//this is the answer

OpenStudy (texaschic101):

2n - 6 = 0 2n = 0 + 6 2n = 6 2n/2 = 6/2 n = 3 check... 3(n + 4) = n + 18 3(3 + 4) = 3 + 18 3(7) = 21 21 = 21 (correct) ANSWER : n = 3